# Fremlin tensor product respects the unbounded order convergence

### v2

@@ Abstract.
Suppose $\Sigma$ is a topological space and $S(\Sigma)$ is the vector lattice of all equivalence classes of continuous real-valued functions defined on open dense subsets of $\Sigma$. In this paper, we establish some lattice and topological aspects of $S(\Sigma)$. In particular, as an application, we show that the unbounded order convergence and the order convergence are stable under passing to the Fremlin tensor product of two Archimedean vector lattices. Furthermore, by considering the Fremlin projective tensor product between Banach lattices, we establish a partial variant of this result for unbounded norm convergence and unbounded absolute weak convergence, as well.
@@ 1. Introduction
Suppose $E$ and $F$ are Archimedean vector lattices. Fremlin in [9] has constructed the Archimedean vector lattice $E\overline{\otimes}F$ (known as the Fremlin tensor product of two Archimedean vector lattices). Under this structure, the algebraic tensor product $E\otimes F$ can be considered as an ordered vector subspace of $E\overline{\otimes}F$. There are various properties of $E$ and $F$ that can be or can not be inherited by the Fremlin tensor product. For example, the Fremlin tensor product behaves well with sublattices; that is if $E_{0}$ and $F_{0}$ are vector sublattices of vector lattices $E$ and $F$, respectively, then $E_{0}\overline{\otimes}F_{0}$ is a vector sublattice of $E\overline{\otimes}F$. Furthermore, if $E_{0}$ and $F_{0}$ are order dense or regular, then, so is the Fremlin tensor product $E_{0}\overline{\otimes}F_{0}$ (for more details, see[13, Proposition 3.2]). However, there are many properties that fail in the Fremlin tensor product between vector lattices even so the underlying spaces possess the corresponding properties, as well. For example, order completeness can not be preserved by the Fremlin tensor product. More precisely, Fremlin in [10, 4C], showed that the Fremlin tensor product of $L^{2}[0,1]$ with itself fails to be order complete even when completed relative to its norm. Also, Grobler in [13, Example 3.8], constructed an example of two compact Hausdorff extremally disconnected topological spaces $K_{1}$ and $K_{2}$ such that $K_{1}\times K_{2}$ is not extremally disconnected so that $C(K_{1}\times K_{2})$ is not order complete, even so both $C(K_{1})$ and $C(K_{2})$ are order complete.
@@
Moreover, the Fremlin tensor product does not respect ideals in general (see [7, Theorem 3.8] for more details). In fact, the authors in [7] have constructed an ideal whose the Fremlin tensor product with itself is not an ideal. Since tensor products are a useful tool in studying other mathematical subjects, it is interesting in its own right, to see whether or not some properties of the underlying spaces can be inherited by the tensor product. One of the possible questions arises while we are dealing with different convergence notions. For example, it is known that the projective tensor norm on the projective tensor product of two Banach spaces is a cross-norm so that it preserves the norm convergence. However, it does not behave well with the weak convergence. More precisely, it is shown in [15, Example 2.10] that the diagonal of the projective tensor product $\ell_{2}\widehat{\otimes}\ell_{2}$ is isometrically isomorphic to $\ell_{1}$. Note that by diagonal, we mean the closed subspace of $\ell_{2}\widehat{\otimes}\ell_{2}$ generated by the elementary tensors $e_{n}\otimes e_{n}$, in which $(e_{n})_{n\in\mathbb{N}}$ is the standard basis of $\ell_{2}$. Moreover, if $u=\Sigma_{n=1}^{\infty}\alpha_{n}e_{n}\otimes e_{n}$, then it can be verified that the projective norm of $u$, denoted by $\pi(u)$, is equal to $\Sigma_{n=1}^{\infty}|\alpha_{n}|$ so that $D$ is isometrically isomorphism to $\ell_{1}$. Now consider the standard basis $(e_{n})\subseteq\ell_{2}$ which is weakly null but $e_{n}\otimes e_{n}$ is not since $\pi(e_{n}\otimes e_{n})=1$ and in $\ell_{1}$, weakly convergence and norm convergence agree. It is known that, in the case of a vector lattice, we have two convergence structures: order convergence and unbounded order convergence. Therefore, it would be interesting to consider the following question:
@@
Does the Fremlin tensor product of two vector lattices preserve either order convergence or unbounded order convergence?
Grobler in [13, Corollary 3.4] considered a ”separate” version answer to this question. More precisely, he proved that for Archimedean vector lattices $E$ and $F$, if $\sigma:E\times F\to E\overline{\otimes}F$ is the natural lattice bimorphism ($(x,y)\rightarrow x\otimes y$), then $\sigma$ is order ($uo$-continuous) on each component. In Section 3, we show that the Fremlin tensor product of two Archimedean vector lattices behave well with both order convergence and unbounded order convergence in a general sense; this extends the result of Grobler to a ”jointly” version, as well. Before we reach to this goal, we need to verify some lattice and topological structures for the vector lattice of all equivalence classes of continuous real-valued functions defined on open dense subsets of a topological space $\Sigma$. Suppose now $E$ and $F$ are Banach lattices. Fremlin in [10] has constructed the Banach lattice $E\widehat{\otimes}F$ (known as the Fremlin projective tensor product of $E$ and $F$); $E\overline{\otimes}F$ is norm dense vector sublattice in $E\widehat{\otimes}F$ so that a normed lattice in its own right. On the other hand, while we are dealing with Banach lattices, we have two known natural notions for unbounded convergences: unbounded norm convergence and unbounded absolute weak convergence. Therefore, the following question is natural.
Question 2.
Does the Fremlin projective tensor product of two Banach lattices preserve either unbounded norm convergence or unbounded absolute weak convergence?
In Section 4, we shall provide a partial positive answer to this question, as well. In Section 2, we provide some preliminaries that are necessary for the main results. In fact, we give a brief exposition on unbounded convergences and also on the Fremlin tensor products, as well.
@@
2.2. the Fremlin tensor products
In this part, we recall some notes about the Fremlin tensor product between vector lattices and also Banach lattices. For more details, see [9, 10]. Furthermore, for a comprehensive, new and interesting reference, see [18]. Furthermore, for a short and nicely written exposition on different types of tensor products between Archimedean vector lattices, see [13].
Suppose $E$ and $F$ are Archimedean vector lattices. In 1972, Fremlin constructed a tensor product $E\overline{\otimes}F$ that is an Archimedean vector lattice with the following properties:
  • The algebraic tensor product $E\otimes F$ is a vector subspace of $E\overline{\otimes}F$ so that it is an ordered vector subspace in its own right.
  • The vector sublattice in $E\overline{\otimes}F$ generated by $E\otimes F$ is the whole of $E\overline{\otimes}F$.
  • For each Archimedean vector lattice $G$ and every lattice bimorphism $\Phi:E\times F\to G$, there is a unique lattice homomorphism $T:E\overline{\otimes}F\to G$ such that $T(x\otimes y)=\Phi(x,y)$ for each $x\in E$ and for each $y\in F$.
Therefore, every element of $E\overline{\otimes}F$ can be considered as a finite suprema and finite infima of some elements of $E\otimes F$. The good news is that we have some density properties for elements of $E\overline{\otimes}F$ in terms of the elements of the algebraic tensor product $E\otimes F$ as follows.
  • Assume that $E$ and $F$ are Archimedean vector lattice. For each $a\in E\overline{\otimes}F$, there exist $u\in E_{+}$ and $v\in F_{+}$ such that for each $\varepsilon>0$, there is $b\in E\otimes F$ with $|a-b|\leq\varepsilon u\otimes v$, [18, Proposition 3.11].
  • For each $c\in(E\overline{\otimes}F)_{+}$, we have $c=\sup\{a\otimes b:a\in E_{+},b\in F_{+}\}$, [18, Proposition 3.12].
  • For each $u\in(E\overline{\otimes}F)_{+}$, there exist $a\in E_{+}$ and $b\in F_{+}$ with $u\leq a\otimes b$, [10, 1A(d)].
Now, suppose $E$ and $F$ are Banach lattices. Fremlin in [10] constructed a tensor product $E\widehat{\otimes}F$ that is a Banach lattice with the following properties.
  • $E\widehat{\otimes}F$ is the norm completion of $E\otimes F$ with respect to the positive-projective norm: for each $u=\Sigma_{i=1}^{n}x_{i}\otimes y_{i}\in E\otimes F$:
[…]
  • $E\overline{\otimes}F$ can be considered as an norm-dense vector sublattice of $E\widehat{\otimes}F$.
  • For every Banach lattice $G$ and every continuous positive bilinear mapping $\Phi:E\times F\to G$, there is a unique positive increasing linear mapping $T:E\widehat{\otimes}F\to G$ with $T(x\otimes y)=\Phi(x,y)$; moreover, this correspodence is norm-preserving.
  • The positive cone in $E\widehat{\otimes}F$ is the closure of the cone $P\subseteq E\otimes F$ generated by $\{x\otimes y:x\in E_{+},y\in F_{+}\}$.
  • The projective norm, .$\|.\|_{\pi}$, on $E\widehat{\otimes}F$ is a cross-norm that is for every $x\in E$ and for every $y\in F$, we have $\|x\otimes y\|_{\pi}=\|x\|\|y\|$.
@@ 3.1. Fremlin tensor product and $uo$-convergence
Suppose $X$ is a topological space and consider the space $C^{\infty}(X)$ consisting of all extended continuous functions from $X$ into the extended reals ($\mathbb{R}^{*}=[-\infty,\infty]$). In general, this space need not be a vector space (see [6, Section 2]). However, when $X$ is extremally disconnected, it is known that $C^{\infty}(X)$ is a vector lattice under the pointwise vector and order operations. Suppose $K_{1}$ and $K_{2}$ are compact Hausdorff spaces. It is known that $C(K_{1})\otimes C(K_{2})$ is norm (order) dense in $C(K_{1}\times K_{2})$. On the other hand, by the known Kakutani’s theorem ([2, Theorem 4.21]), every Archimedean vector lattice with an order unit can be considered as a norm (order) vector sublattice of some $C(K)$-space ($K$ compact and Hausdorff). So, we can transfer problems regarding tensor products of Archimedean vector lattices with order unit in terms of $C(K)$-spaces. However, the issue is that Archimedean vector lattices with order units are scarce in the category of all vector lattices. On the other hand, by the Maeda-Ogasawara theorem ([1, Theorem 7.29]), every Archimedean vector lattice can be considered as an order dense vector sublattice of some $C^{\infty}(X)$-space for some compact Hausdorff extremally disconnected topological space $X$. Now, if we want to develop a similar theory for the tensor product between arbitrary Archimedean vector lattices in terms of $C^{\infty}(X)$-spaces, the issue is that the Cartesian product of two extremally disconnected topological spaces $X$ and $Y$, is not extremally disconnected, in general. So, $C^{\infty}(X\times Y)$ may not a vector lattice (see [13, Example 3.8]). This problem was remarkably solved by Buskes and Wickstead in 2017 ([6]) by introducing a new larger vector lattice.
@@
Suppose $\Sigma$ is a topological space. By $S(\Sigma)$, we mean the space of all equivalence classes of continuous functions defined on open dense subsets of $\Sigma$ under the equivalence relation $f\thicksim g$ if they coincide on the intersection of their domains. This space is introduced by Buskes and Wickstead in [6]. It is a vector lattice under the pointwise lattice and vector operations.
Note that elements of a $C^{\infty}(X)$-space ($X$ a topological space) are functions, nevertheless, while we are working with a $S(\Sigma)$-space ($\Sigma$ a topological space), we are dealing with some classes of functions in which the class of each function, contains all of functions that are equal on the intersection of their domains. This defines an equivalence relation on $S(\Sigma)$. The proof of the following lemma follows from the equivalence relation defined on $S(\Sigma)$.
Lemma 1.
Suppose $\Sigma$ is a topological space and $a\in S(\Sigma)$. If $f,g\in a$ and $t\in Dom(f)\cap Dom(g)$ for some $t\in\Sigma$, then $f(t)=g(t)$.
Now, we consider a convention: for each $a\in S(\Sigma)$, put ${Dom}(a):=\cup_{f\in a}{\text{D}om}(f)$; this is well-defined. By $a(t)=m$ also we mean there exists $f\in a$ such that $f(t)=m$. By [18, Lemma 2.2], we see that for each $f\in S(\Sigma)$, there exists an element in its class with the maximal domain so that when picking a representation for an equivalence class in $S(\Sigma)$, we may assume that it is the one with maximal domain. Therefore, for a continuous real-valued function $f$ defined on an open dense subset $\Omega_{f}$ of $\Sigma$, we can identify $f$ with its equivalence class $[f]$. With this convention, assume that $X$ is an extremally disconnected topological space. Then by restriction of every $f\in C^{\infty}(X)$ to the preimage of the reals which is an open dense subset of $X$, ([1, Definition 7.26]), we see that $C^{\infty}(\Sigma)\subseteq S(\Sigma)$. Moreover, every continuous function $f\in C(\Sigma)$ can be identified with $[f]\in S(\Sigma)$, in which, $[f]$ is the class consisting of all $g\in S(\Sigma)$ with $f\thicksim g$. Therefore, we can have the inclusion $C(\Sigma)\subseteq S(\Sigma)$.
When $\Sigma$ is extremally disconnected, there is a nice relation between $S(\Sigma)$ and $C^{\infty}(\Sigma)$.
@@
Proof.
By [1, Lemma 7.25], for each $f\in S(\Sigma)$, there exists a unique extension $f^{*}\in C^{\infty}(\Sigma)$. Define $\Phi:S(\Sigma)\to C^{\infty}(\Sigma)$ defined via $\Phi(f)=f^{*}$. We show that $\Phi$ is a lattice isomorphism. First, note that $\Phi$ is well-defined as the extension is unique. $\Phi$ is linear. Suppose $f:\Omega_{1}\to\mathbb{R}$ and $g:\Omega_{2}\to\mathbb{R}$ are elements of $S(\Sigma)$. Put $\Omega=\Omega_{1}\cap\Omega_{2}$ which is an open dense subset of $\Sigma$, again. Note that $f^{*}+g^{*}$ is an extension of $f+g$ defined on $\Omega$. So, by uniqueness of the extension, we have $\Phi(f+g)=(f+g)^{*}=f^{*}+g^{*}=\Phi(f)+\Phi(g)$. Now, assume that $h\in C^{\infty}(\Sigma)$. By the definition of the extended functions, $\Omega_{h}=h^{-1}(\mathbb{R})=\{x\in\Sigma,-\infty<h(x)<\infty\}$ is an open dense subset of $\Sigma$ so that the restriction of $h$ to $\Omega_{h}$ (denoted by $h_{0}$) lies in $S(\Sigma)$. It is easy to see that $\Phi(h_{0})=h$. Therefore, $\Phi$ is onto. Suppose, $\Phi(f)=f^{*}=0$ so that for each $t\in\Omega_{1}$, we have $f(t)=0$. Therefore, $\Phi$ is also one-to-one. Moreover, for each $t\in\Omega$, we have, $(f\vee g)^{*}(t)=(f\vee g)(t)=f(t)\vee g(t)=f^{*}(t)\vee g^{*}(t)$. So, again, uniqueness of the extension, results in $\Phi(f\vee g)=(f\vee g)^{*}=f^{*}\vee g^{*}=\Phi(f)\vee\Phi(g)$. This completes the proof. ∎
@@
We have seen that $C(\Sigma)$, the space of all real-valued continuous functions on topological space $\Sigma$, can be considered as a vector sublattice of $S(\Sigma)$. But, we have more if we consider completely regular topological spaces.
@@
Now, we characterize $uo$-convergence in $S(\Sigma)$. The proof is similar to the proof of [4, Theorem 7.1]; see also [5, Theorem 3.2].
@@
Suppose $\Sigma$ is a completely regular topological space. For a net $(f_{\alpha})\subseteq{S(\Sigma)}_{+}$, $f_{\alpha}\xrightarrow{uo}0$ if and only if for each non-empty open set $U\subseteq\Sigma$ and for each $\varepsilon>0$, there exist a non-empty open set $V\subseteq U$ and an index $\alpha_{0}$ such that $f_{\alpha}(t)\leq\varepsilon$ for each $t\in V\cap Dom(f_{\alpha})$.
@@
Suppose $f_{\alpha}\xrightarrow{uo}0$. Assume that $U\subseteq\Sigma$ is a non-empty open set and $\varepsilon>0$ is arbitrary. Since $\Sigma$ is completely regular, there exists a non-zero positive continuous function $g$ on $\Sigma$ with $g\leq\textbf{1}_{U}$. By [4, Theorem 6.2], we can find non-zero positive function $h\in S(\Sigma)$ and an $\alpha_{0}$ such that $(f_{\alpha}-\varepsilon h)^{+}\perp g$ for each $\alpha\geq\alpha_{0}$. Therefore, $(f_{\alpha}-\varepsilon h)^{+}$ vanishes on non-empty open subset $V:=supp(g)\subseteq U$. Thus, for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$, $f_{\alpha}(x)\leq\varepsilon$.
For the other direction, Take any non-zero positive $h\in S(\Sigma)$. There are non-empty open set $U\subseteq\Sigma$ and $\varepsilon>0$ with $h\geq\varepsilon\textbf{1}_{U}$. By the assumption, There are non-empty open set $V\subseteq U$ and an $\alpha_{0}$ such that $f_{\alpha}(x)\leq\varepsilon$ for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$. Since $\Sigma$ is completely regular, there is a non-zero positive continuous function $g$ with $g\leq\varepsilon\textbf{1}_{U}$. Now, we can see that $(f_{\alpha}-h)^{+}$ vanishes on $V$ so that disjoint with $g$. Again, using [4, Theorem 6.2], convinces us that $f_{\alpha}\xrightarrow{uo}0$. ∎
@@
Suppose $h\in{S(\Sigma\times\Omega)}_{+}$. There exists $(t_{0},s_{0})\in\Sigma\times\Omega$ such that $h(t_{0},s_{0})=r>0$. Since $h$ is continuous, there exist non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $h(t,s)>\frac{r}{2}$ for all $(t,s)\in U\times V$. Since $\Sigma$ and $\Omega$ are completely regular, we can find $0\neq f\in{S(\Sigma)}_{+}$ such that $f\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Sigma)}$ and $f$ vanishes outside of $U$. Also, there is $0\neq g\in{S(\Omega)}_{+}$ with $g\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Omega)}$ and $g=0$ outside of $V$. We show that $0<f\otimes g\leq h$. Since both $f$ and $g$ are non-zero, $f\otimes g>0$. Suppose $(t,s)\in(Dom(f)\times Dom(g))\cap Dom(h)$. First assume that $(t,s)\in U\times V$. Then, $(f\otimes g)(t,s)=f(t)g(s)\leq\frac{r}{2}<h(t,s)$. If $(t,s)\notin U\times V$, then either $f(t)=0$ or $g(s)=0$. Therefore, $(f\otimes g)(t,s)=f(t)g(s)=0\leq h(t,s)$. ∎
The proof of the following result follows from [13, Proposition 3.2].
@@
Suppose $E$ and $F$ are Archimedean vector lattices. Moreover, assume that $f_{\alpha}\xrightarrow{uo}f$ in $E$ and $g_{\alpha}\xrightarrow{uo}g$ in $F$. Then, $f_{\alpha}\otimes g_{\alpha}\rightarrow f\otimes g$ in the Fremlin tensor product $E\overline{\otimes}F$.

### v4

# Fremlin tensor product behaves well with the unbounded order convergence

@@ Abstract.
Suppose $\Sigma$ is a topological space and $S(\Sigma)$ is the vector lattice of all equivalent classes of continuous real-valued functions defined on open dense subsets of $\Sigma$. In this paper, we establish some lattice and topological aspects of $S(\Sigma)$. In particular, as an application, we show that the unbounded order convergence and the order convergence are stable under passing to the Fremlin tensor product of two Archimedean vector lattices.
@@ 1. Introduction
Suppose $E$ and $F$ are Archimedean vector lattices. Fremlin in [7] has constructed the Archimedean vector lattice $E\overline{\otimes}F$ (known as the Fremlin tensor product of two Archimedean vector lattices). Under this structure, the algebraic tensor product $E\otimes F$ can be considered as an ordered vector subspace of $E\overline{\otimes}F$ with some valuable density properties (see [8, 1A] for details). There are various properties of $E$ and $F$ that can be or can not be inherited by the Fremlin tensor product. For example, the Fremlin tensor product behaves well with sublattices; that is if $E_{0}$ and $F_{0}$ are vector sublattices of $E$ and $F$, respectively, then $E_{0}\overline{\otimes}F_{0}$ is a vector sublattice of $E\overline{\otimes}F$. Furthermore, if $E_{0}$ and $F_{0}$ are order dense or regular, then, so is the Fremlin tensor product $E_{0}\overline{\otimes}F_{0}$ (for more details, see[10, Proposition 3.2]). However, there are many properties that fail in the Fremlin tensor product even for some classical spaces. For example, order completeness can not be preserved by the Fremlin tensor product ([10, Example 3.8] or [8, Example 4C]). Moreover, the Fremlin tensor product does not respect ideals or bands in general (see [6] for more details). Since tensor products are a useful tool in studying other mathematical subjects, it is interesting in its own right, to see whether or not some properties of the underlying spaces can be inherited by the tensor product. One of the possible questions arises while we are dealing with different convergence notions. For example, it is known that the projective tensor norm on the projective tensor product of two Banach spaces is a cross-norm so that it preserves the norm convergence. However, it does not respect the weak convergence; it is shown in [11, Example 2.10] that the diagonal of the projective tensor product $\ell_{2}\widehat{\otimes}\ell_{2}$ is isometrically isomorphic to $\ell_{1}$. Note that by the diagonal, we mean the closed subspace of $\ell_{2}\widehat{\otimes}\ell_{2}$ generated by the elementary tensors $e_{n}\otimes e_{n}$, in which $(e_{n})_{n\in\mathbb{N}}$ is the standard basis of $\ell_{2}$. To see this, observe that if $u=\Sigma_{n=1}^{\infty}\alpha_{n}e_{n}\otimes e_{n}$, then it can be verified that the projective norm $u$, denoted by $\pi(u)$, is equal to $\Sigma_{n=1}^{\infty}|\alpha_{n}|$ so that $D$ is isometrically isomorphism to $\ell_{1}$. Now consider the standard basis $(e_{n})\subseteq\ell_{2}$ which is weakly null but $e_{n}\otimes e_{n}$ is not since $\pi(e_{n}\otimes e_{n})=1$ and in $\ell_{1}$, weakly convergence and norm convergence agree. It is known that, in the case of a vector lattice, we have two convergence structures: order convergence and unbounded order convergence. Therefore, it would be interesting to consider the following question:
@@
Does the Fremlin tensor product of two vector lattices respect either order convergence or unbounded order convergence?
Grobler in [10, Corollary 3.4] considered a ”separate” version answer to this question. More precisely, he proved that for Archimedean vector lattices $E$ and $F$, if $\sigma:E\times F\to E\overline{\otimes}F$ is the natural lattice bimorphism ($(x,y)\rightarrow x\otimes y$), then $\sigma$ is order ($uo$-continuous) on each component. In this note, we show that the Fremlin tensor product of two Archimedean vector lattices behaves well with both order convergence and unbounded order convergence in a general sense; this extends the result of Grobler to a ”jointly” version, as well. Before we reach to this goal, we need to verify some lattice and topological structures for the vector lattice consisting of equivalent classes of all continuous real-valued functions defined on open dense subsets of a topological space $\Sigma$.
@@
In this part, we recall some notes about the Fremlin tensor product between vector lattices. For more details, see [7, 8]. Furthermore, for a comprehensive, new and interesting reference, see [Wickstead1:24]. Furthermore, for a short and nicely written exposition on different types of tensor products between Archimedean vector lattices, see [10].
Suppose $E$ and $F$ are Archimedean vector lattices. In 1972, Fremlin constructed a tensor product $E\overline{\otimes}F$ that is an Archimedean vector lattice with the following features:
  • The algebraic tensor product $E\otimes F$ is a vector subspace of $E\overline{\otimes}F$. Therefore, we conclude that it is an ordered vector space in its own right.
  • The vector sublattice in $E\overline{\otimes}F$ generated by the algebraic tensor product $E\otimes F$ is $E\overline{\otimes}F$.
  • For each Archimedean vector lattice $G$ and every lattice bimorphism $\Phi:E\times F\to G$, there is a unique lattice homomorphism $T:E\overline{\otimes}F\to G$ such that $T(x\otimes y)=\Phi(x,y)$ for all $x\in E$ and for all $y\in F$.
Therefore, we can consider every element of $E\overline{\otimes}F$ as a finite supremum and finite infimum of some elements of $E\otimes F$. Now, assume that $E$ is a vector lattice. It is shown in [12] that $E$ has a representation as a vector sublattice of some $S(\Sigma)$-space, where $\Sigma$ is a topological space and $S(\Sigma)$ is the vector lattice of all (equivalence classes) of continuous real-valued functions defined on open dense subsets of $\Sigma$ (we shall speak about this space more in the next section). Now, we recall [5, Proposition 3.1] that is crucial for the rest of the paper.
Assume that $E$ and $F$ are vector lattices with the representations as vector sublattices of some $S(\Sigma)$ and $S(\Omega)$ spaces, respectively, where $\Sigma$ and $\Omega$ are topological spaces. Then the Fremlin tensor product $E\overline{\otimes}F$ is isomorphic to the vector sublattice of $S(\Sigma\times\Omega)$ generated by the functions $(\sigma,\gamma)\rightarrow x(\sigma)y(\gamma)$, for each $x\in E$ and for each $y\in F$.
@@
Suppose $X$ is a topological space and consider the space $C^{\infty}(X)$ consisting of all continuous extended real-valued functions that are finite except on a nowhere dense set. In general, this space need not be a vector space (see [5, Section 2]). However, when $X$ is extremally disconnected, it is known that $C^{\infty}(X)$ is a vector lattice under the pointwise vector and order operations. Suppose $K_{1}$ and $K_{2}$ are compact Hausdorff spaces. It is known that $C(K_{1})\otimes C(K_{2})$ is norm (order) dense in $C(K_{1}\times K_{2})$. On the other hand, by the known Kakutani’s theorem ([2, Theorem 4.21]), every Archimedean vector lattice with an order unit can be considered as an norm (order) vector sublattice of some $C(K)$-space ($K$ compact and Hausdorff). So, we can transfer problems regarding tensor product of Archimedean vector lattices with order units in terms of $C(K)$-spaces. However, the issue is that Archimedean vector lattices with order units are scarce in the category of all vector lattices. On the other hand, by the Maeda-Ogasawara theorem ([1, Theorem 7.29]), every Archimedean vector lattice can be considered as an order dense vector sublattice of some $C^{\infty}(X)$-space for some compact Hausdorff extremally disconnected topological space $X$. Now, if we want to develop a theory for tensor products in terms of $C^{\infty}(X)$-spaces, the problem is that the Cartesian product of two extremally disconnected topological spaces $X$ and $Y$, is not extremally disconnected, in general. So, $C^{\infty}(X\times Y)$ may not a vector lattice (see [10, Example 3.8]). This problem was remarkably solved by Buskes and Wickstead in 2017 ([5]) by introducing a new larger vector lattice.
@@
Suppose $\Sigma$ is a topological space. By $S(\Sigma)$, we mean the space of all equivalence classes of continuous functions defined on open dense subsets of $\Sigma$ under the equivalence relation $f\thicksim g$ if they coincide on the intersection of their domains. This space is introduced by Buskes and Wickstead in [5]. It is a vector lattice under the pointwise lattice and vector operations. Observe that any $f\in S(\Sigma)$ is in fact a class of functions; more precisely, $[f]=\{g\in S(\Sigma),f\thicksim g\}$. Now, define $\phi$ from $C^{\infty}(\Sigma)$ into $S(\Sigma)$ defined via $\Phi(f)=[f]$; note that, in this case, we restrict every $f\in C^{\infty}(\Sigma)$ to the open dense subset whose image is real. $\Phi$ is clearly injective so that we may identify $C^{\infty}(\Sigma)$ with its image under $\Phi$ to think $C^{\infty}(\Sigma)$ as a subset of $S(\Sigma)$. In particular, $C(\Sigma)$ can be treated as a subset (in fact, a vector sublattice) of $S(\Sigma)$.
Now, assume that $\Sigma$ is also extremally disconnected. Then $\Phi$ is a vector lattice homomorphism by the definition of vector and lattice operations in $S(\Sigma)$. Furthermore, $\Phi$ is onto by using [1, Theorem 7.25]. So, we have the following observation.
@@
Note that $C(\Sigma)$, the space of all real-valued continuous functions on topological space $\Sigma$, can be considered as a vector sublattice of $S(\Sigma)$. But, we have more if we consider completely regular topological spaces.
@@
Now, we characterize $uo$-convergence in $S(\Sigma)$. The proof is similar to the proof of [3, Theorem 7.1]; see also [4, Theorem 3.2].
@@
Suppose $\Sigma$ is a completely regular topological space. For a net $(f_{\alpha})\subseteq{S(\Sigma)}$, $f_{\alpha}\xrightarrow{uo}0$ if and only if for each non-empty open set $U\subseteq\Sigma$ and for each $\varepsilon>0$, there exist a non-empty open set $V\subseteq U$ and an index $\alpha_{0}$ such that $|f_{\alpha}(t)|\leq\varepsilon$ for each $t\in V\cap Dom(f_{\alpha})$.
@@
Suppose $f_{\alpha}\xrightarrow{uo}0$. Suppose $U\subseteq\Sigma$ is a non-empty open set and $\varepsilon>0$ is arbitrary. Since $\Sigma$ is completely regular, there exists a non-zero positive continuous function $g$ on $\Sigma$ with $g\leq\textbf{1}_{U}$. By [3, Theorem 6.2], we can find non-zero positive function $h\in S(\Sigma)$ and an $\alpha_{0}$ such that $(|f_{\alpha}|-\varepsilon h)^{+}\perp g$ for each $\alpha\geq\alpha_{0}$. Therefore, $(|f_{\alpha}|-\varepsilon h)^{+}$ vanishes on non-empty open subset $V:=supp(g)\subseteq U$. Thus, for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$, $|f_{\alpha}(x)|\leq\varepsilon$.
For the other direction, take any non-zero positive $h\in S(\Sigma)$. There are non-empty open set $U\subseteq\Sigma$ and $\varepsilon>0$ with $h\geq\varepsilon\textbf{1}_{U}$. By the assumption, there are non-empty open set $V\subseteq U$ and an $\alpha_{0}$ such that $|f_{\alpha}(x)|\leq\varepsilon$ for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$. Since $\Sigma$ is completely regular, there is a non-zero positive continuous function $g$ with $g\leq\varepsilon\textbf{1}_{U}$. Now, we can see that $(|f_{\alpha}|-h)^{+}$ vanishes on $V$ so that disjoint with $g$. Again, using [3, Theorem 6.2], convinces us that $f_{\alpha}\xrightarrow{uo}0$. ∎
@@
Suppose $h\in{S(\Sigma\times\Omega)}_{+}$. There exists $(t_{0},s_{0})\in\Sigma\times\Omega$ such that $h(t_{0},s_{0})=r>0$. Since $h$ is continuous, there exist non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $h(t,s)>\frac{r}{2}$ for all $(t,s)\in U\times V$. Since $\Sigma$ and $\Omega$ are completely regular, we can find $0\neq f\in{S(\Sigma)}_{+}$ such that $f\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Sigma)}$ and $f$ vanishes outside of $U$. Also, there is $0\neq g\in{S(\Omega)}_{+}$ with $g\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Omega)}$ and $g=0$ outside of $V$. We show that $0<f\otimes g\leq h$. Since both $f$ and $g$ are non-zero, $f\otimes g>0$. Suppose $(t,s)\in Dom(f)\cap Dom(g)\cap Dom(h)$. First assume that $(t,s)\in U\times V$. Then, $(f\otimes g)(t,s)=f(t)g(s)\leq\frac{r}{2}<h(t,s)$. If $(t,s)\notin U\times V$, then either $f(t)=0$ or $g(s)=0$. Therefore, $(f\otimes g)(t,s)=f(t)g(s)=0\leq h(t,s)$. ∎
The proof of the following result follows from [10, Proposition 3.2].
@@
Suppose $E$ and $F$ are Archimedean vector lattices. Moreover, assume that $f_{\alpha}\xrightarrow{uo}f$ in $E$ and $g_{\beta}\xrightarrow{uo}g$ in $F$. Then, $f_{\alpha}\otimes g_{\beta}\rightarrow f\otimes g$ in the Fremlin tensor product $E\overline{\otimes}F$.